# InternLM2.5-StepProver: Advancing Automated Theorem Proving via Expert Iteration on Large-Scale LEAN Problems

### v1

@@ Abstract
Large Language Models (LLMs) have emerged as powerful tools in mathematical theorem proving, particularly when utilizing formal languages such as LEAN. The major learning paradigm is expert iteration, which necessitates a pre-defined dataset comprising numerous mathematical problems. In this process, LLMs attempt to prove problems within the dataset and iteratively refine their capabilities through self-training on the proofs they discover. We propose to use large-scale LEAN problem datasets Lean-workbook for expert iteration with more than 20,000 CPU days. During expert iteration, we found log-linear trends between solved problem amount with proof length and CPU usage. We train a critic model to select relatively easy problems for policy models to make trials and guide the model to search for deeper proofs. InternLM2.5-StepProver achieves open-source state-of-the-art on MiniF2F, Lean-Workbook-Plus, ProofNet, and Putnam benchmarks. Specifically, it achieves a pass of 65.9% on the MiniF2F-test and proves (or disproves) 17.0% of problems in Lean-Workbook-Plus which shows a significant improvement compared to only 9.5% of problems proved when Lean-Workbook-Plus was released. We open-source our models and searched proofs at https://github.com/InternLM/InternLM-Math and https://huggingface.co/datasets/internlm/Lean-Workbook.
@@ 1 Introduction
Automated theorem proving has been a challenging topic in artificial intelligence (Pfenning, 2004; Zheng et al., 2021; Wu et al., 2022; Polu et al., 2022) which requires complex reasoning and a deep understanding of math. AlphaProof [^1] https://deepmind.google/discover/blog/ai-solves-imo-problems-at-silver-medal-level/ have demonstrated remarkable progress by achieving silver-medal performance on International Mathematical Olympiad problems using the LEAN 4 proof assistant, particularly excelling in number theory and algebra. AlphaProof’s training regime, based on AlphaZero methodology (Silver et al., 2017), encompasses 100 million formal mathematics problems—a scale that significantly surpasses previous efforts (Polu et al., 2022; Lample et al., 2022; Xin et al., 2024a; b). Building on this advancement, our work leverages the Lean-workbook (Ying et al., 2024a), the largest open-source problem collection available, to conduct systematic expert iteration and analyze proving strategies at scale.
Our extensive experimentation, consuming over 20,000 CPU days, yielded several key insights into the challenge of automated theorem proving:
  • Success Rate: Only 1.5% of CPU usage resulted in successful proofs or disproofs, highlighting the substantial difficulty of automated theorem proving.
  • Search Strategy: The integration of a critic model consistently produced deeper proofs compared to naive best-first-search approaches.
@@
  • Scaling Relationships: We identified a log-linear relationship between the number of proved problems and both proof length and computational resources, providing valuable insights for resource allocation in future work.
These findings offer practical guidance for enhancing proof search efficiency and suggest pathways toward scaling to larger training sets comparable to AlphaProof’s 100 million problems. We evaluated our final models against established benchmarks, including MiniF2F (Zheng et al., 2021), Putnam (Tsoukalas et al., 2024), and ProofNet (Azerbayev et al., 2023a). Our results demonstrate state-of-the-art performance among open-source systems, achieving better results to Deepseek-Prover-V1.5 (Xin et al., 2024b).
@@ 2 Methods
Based on Lean-workbook-plus Ying et al. (2024a), one of the largest auto-formalized problem sets in Lean 4, we propose to exploit the potential of formal reasoning by expert iteration (Anthony et al., 2017; Polu et al., 2022).
Expert Iteration on Lean-Workbook-Plus
@@
For each statement in Lean-Workbook-plus, we try to prove or disprove it following Xin et al. (2024a). We search proofs based on best-first-search and critic-guided search via generating a tactic as an action (Polu et al., 2022; Azerbayev et al., 2023b; Wu et al., 2024). Starting with InternLM2-StepProver (Wu et al., 2024), our latest model for formal reasoning, we have eliminated the cumbersome bootstrap process. Initially, we conduct a rapid scan of the entire Lean-workbook-plus dataset using a relatively small search budget (i.e. 10 iterations at most per problem and a time limit of 50 seconds). The discovered proofs are added to the training set, and the solved problems, along with their negated statements, are removed from the dataset. This process helps us identify statements that are inherently unprovable, thereby enhancing the efficiency of the iteration.
@@
We then repeat this process in multiple rounds, gradually increasing the search budget for subsequent evaluations until a predefined upper bound (at most 2000 iterations and 3600 seconds per problem) is reached. After each round, we retrain our policy and critic models using an expanded set of successful proof trajectories. Since some founded proofs are ill-formed and contain many irrelevant proof steps with larger CPU consumptions. We continue to search for proofs for these problems and use discovered shorter and more learnable proofs to improve our models. After model iterations, we use the critic model to re-evaluate all unproved statements with scores and we only focus on search proofs on the top 50% of problems that the model is most likely to be solvable.
Updating Policy Model
@@
Updating Critic Model
From our observation, using best-first-search with log-probability scores seldom leads to deep proofs (shown in Figure 1), and limits the proof ability of our model. Therefore, we decide to train a critic model Lample et al. (2022); Polu et al. (2022) to better guide our policy model for proof generation. The critic model ($V$) uses the proof state ($s$) as the input and outputs a scalar ($V(s)\in\mathbb{R}$). We train our critic model in a preference style which is similar to reward model training in RLHF (Ouyang et al., 2022) instead of binary targets (the state can be proved or not). We create two types of preference pairs in our training:
  • Path Pairs: For a successful proof path from the root (initial proof state) to no_goals, we create pairs where the positive example is a state closer to no_goals and the negative example is a state closer to the root (i.e. $V(s_{t})<V(s_{t+\Delta})$). This design implicitly states that any legal tactic leads to no_goals having a positive reward $r$. For a path of length n, this methodology allows us to generate at most ($n\choose 2$ pairs of positive and negative examples.
[…]
Figure 1: Critic-guided search finds more deep proofs compared to best-first-search in ProofNet and MiniF2F. We calculate the proof length based on the line count of the shortest proof for one problem.
@@ 3.1 Results of Expert Iteration
As Tab. LABEL:tab:leanwkbk_results shows, we are able to prove/disprove a total of 17.0% of the Lean-workbook-plus problems, making a noticeable improvement from we release Lean-workbook-plus. These proved and disproved statements and corresponding tactics and states have been released at https://huggingface.co/datasets/internlm/Lean-Workbook. We also unveiled more facts about the expert iteration process, including the distribution of the auto-formalized problem set, the efficacy of our method, etc.
@@
The consumption of CPU/GPU resources.
The search process involves a collaboration of CPU and GPU resources. Given the fixed amount of active GPUs and CPUs, the GPU time consumed is proportional to the total CPU time (in our case, approximately 1:11). In summary, we consumed approximately 21,364 CPU days throughout the entire expert iteration process. However, these search budgets are not uniformly distributed across all formalized problems. Easier problems are more likely to be solved in the early rounds of iteration, thereby ceasing to consume search resources in later rounds. The search consumption of each problem is a key indicator of the distribution of problem difficulty and can provide valuable insights for further scaling. In our case, we selected the CPU time consumed per successful proof as an estimate of resource consumption expectations.Consider the set of problems $S=\{s_{i}\}$. For each problem $s_{i}$, let $P_{s_{i}}$ denote the set of all attempts, and $T_{s_{i},j}$ represent the time spent on the $j$-th attempt of problem $s_{i}$. Define the indicator function $\text{valid}(s_{i},j)$, which equals 1 if the $j$-th attempt on problem $s_{i}$ results in a valid solution, and 0 otherwise. The CPU time consumed per successful proof, $C_{s_{i}}$, is given by: $C_{s_{i}}=\frac{\sum_{j}T_{s_{i},j}}{\sum_{j}\text{valid}(s_{i},j)}$. Table 2 presents a detailed analysis of computational resource consumption. A key observation is that the majority of CPU resources are expended on problems that are challenging to prove or disprove. Notably, only about 1.5% of CPU resources are used to solve 17.0% of the problems, while the remaining 98.5% of resources yield no successful outcomes. Figure 2 provides a more granular analysis of the distribution of CPU search time for each problem. The graph reveals a peak in the near-zero region, suggesting the presence of numerous trivial problems in our dataset. Additionally, a log-linear trend is evident for problems with CPU search times between approximately 10,000 and 30,000 seconds, indicating potential areas for further investigation.
@@
A similar trend is observed within the class of proofs with the same proof lengths. As shown in Fig. 4, when grouping proofs with the same lengths, an approximate log-linear decrease is consistently noticeable. As the proof length increases, the average CPU time also increases, and a delayed peak can be observed. However, even the shortest proofs can be challenging to find. This may be due to the growth in the model’s reasoning ability during iterations.
Figure 2: The distribution of CPU search time per successful proof, which is defined as the total CPU resources spent on searching the specific problem divided by the number of independent successful search trials. Only problems solved are represented in the figure. 83% of the Lean-workbook problems that remain unproven are not depicted.
The proof lengths’ distribution.
Vanilla best-first-search (BF) methods suffer from an explosion of the search space, making it impossible to find long proofs. In contrast, critique-guided (CG) methods are capable of finding longer solutions. The average length of solutions found by the BF method is 1.66, whereas the same indicator is 4.44 for the CG method, demonstrating a significant improvement in deeper reasoning abilities. The distribution of proofs during our expert iteration is shown in Fig. 4. To avoid the impact of redundancies, we only consider the shortest proof for each problem in our analysis. We observe a nearly log-linear trend in the distribution of proof lengths. We state that the unproved 83% problems can have longer proof lengths and require much more time to find them based on our findings.
Figure 3: The distribution of CPU search time with regard to shortest proof length.

### v2

# InternLM2.5-StepProver: Advancing Automated Theorem Proving via Critic-Guided Search

@@ Abstract
Large Language Models (LLMs) have emerged as powerful tools in mathematical theorem proving, particularly when utilizing formal languages such as LEAN. A prevalent proof method involves the LLM prover iteratively constructing the proof tactic by tactic, typically following a best-first search scheme. However, this method often ignores the critical preference information inside the existing tactic trajectories, hindering the search for deeper proofs. We propose an intuitive yet effective method, which utilizes a critic model to capture the preference information and to guide the search of the prover model at runtime. Given the prover-critic framework, a large-scale expert iteration with more than 20,000 CPU days is then applied to further fine-tune the prover and the critic. The trained InternLM2.5-StepProver critic significantly boosts the performance of the prover model ($59.4\%\to 65.9\%$). We also analyze the impact of the critic on various aspects of the theorem proving process during expert iteration, providing insights into its effectiveness. The models and the discovered proofs will be open-sourced.
Machine Learning, Automated Theorem Proving, LEAN, Critic-Guided Search
@@ 1 Introduction
Automated theorem proving has been a challenging topic in artificial intelligence (Pfenning, 2004; Zheng et al., 2021; Wu et al., 2022; Polu et al., 2022) which requires complex reasoning and a deep understanding of mathematics. AlphaProof [^1] https://deepmind.google/discover/blog/ai-solves-imo-problems-at-silver-medal-level/ has demonstrated remarkable progress by achieving silver-medal performance on International Mathematical Olympiad problems using the LEAN 4 proof assistant, particularly excelling in number theory and algebra. The training regime of AlphaProof, based on the AlphaZero methodology (Silver et al., 2017), encompasses 100 million formal mathematics problems—a scale that significantly surpasses previous efforts (Polu et al., 2022; Lample et al., 2022; Xin et al., 2024a, b).
Existing open-source methods typically train a language model on tuples of (proof state, next tactic), followed by best-first tree search to find proofs (Polu & Sutskever, 2020; Polu et al., 2022; Lample et al., 2022; Yang et al., 2024; Lin et al., 2024; Wu et al., 2024). The term ”best-first” refers to relying on internal indicators from the prover model (e.g., the log-probability scores of generated tactics) to guide the search. An observation is that such indicators become unreliable when searching for deeper proofs, leading to performance degradation. Experimental results from previous studies (Wu et al., 2024) have shown that even when the prover model is trained on extensive formal language corpora, the ability to discover deeper proofs remains difficult to acquire. For example, although InternLM2-StepProver achieves 48.8% on miniF2F-test, the longest proof it discovers contains only 8 tactics—a stark contrast to typical tree searches that require 100 to 1000 steps.
Inspired by recent advances in informal mathematical reasoning (Lightman et al., 2023; Xu et al., 2024), our key insight is that the tree structure constructed during the search process is naturally suited for process supervision. The search trajectories inherently provide the data needed to train the supervision model—which we term the critic model. Therefore, a prover-critic framework is built in which the critic model estimates the ”value” of each state and decides which states to explore further.
Building on this framework, our work then leverages the Lean-workbook (Ying et al., 2024a), the largest open-source problem collection available, to conduct systematic expert iteration and analyze proving strategies at scale. We initiate InternLM2.5-StepProver-critic by collecting roughly 8000 preference pairs from the search trajectories when InternLM2-StepProver is evaluated on its training sets. We then perform expert iteration on Lean-workbook. Our extensive experimentation, consuming over 20,000 CPU days, yielded several key contributions to the challenge of automated theorem proving:
  • We present a novel approach that significantly enhances theorem proving capabilities by utilizing a separate critic model to guide the search for deeper proofs. Experiments demonstrate that the critic model boosts the performance of the prover model from $59.4\%$ to $65.6\%$ on miniF2F (when used without the vanilla best-first search method) or to $65.9\%$ when combined with it.
  • We provide an example of massive expert iteration, demonstrating the effectiveness of expert iteration on synthetic datasets. The enlarged Lean-workbook solution set will be open-sourced.
@@
  • We identify a log-linear relationship between the number of proved problems and both proof length and computational resources, providing valuable insights for resource allocation in future work.
Figure 1: Comparison of Best-first Search and Critic-Guided Search methods. Left: Best-first search relies solely on log probability scores from tactic logits, under-utilizing intermediate state information. Right: Critic-Guided search employs a critic model to re-rank states, enabling more informed state expansion decisions. Note that a lower log probability score indicates a higher priority, whereas a higher critique score also signifies increased priority.
@@ 2 Methods
We introduce critic-guided search, a novel framework that enhances automated theorem proving by incorporating critical evaluation mechanisms into the proof search process, to address key limitations of existing best-first search methods through structured critique generation and expert iteration (Anthony et al., 2017; Polu et al., 2022). The methodology consists of two main components: First, we present a critic-guided search algorithm and compare it with traditional best-first search approaches (§2.1). Second, we describe our expert iteration framework that iteratively refines the proof search strategy through bootstrapping processes.
2.1 Search Methods
Best-First Search.
The most widely adopted approach for evaluating the theorem proving capabilities of language models $M$ involves employing best-first search algorithms, as demonstrated in GPT-f (Polu & Sutskever, 2020; Azerbayev et al., 2023b). This method maintains a collection of all unexpanded states $s_{i}$. At each iteration, the algorithm selects the optimal state $s_{i}$ for expansion and employs the language model to generate $S$ candidate tactics $t_{i,1\cdots S}$ for the current state $s_{i}$. For each valid tactic $t_{i,j}$, a new state is derived by executing tactic $t_{i,j}$ on state $s_{i}$.
Following established conventions (Polu & Sutskever, 2020; Yang et al., 2024), the state with the highest negative log-probability is considered optimal. Specifically, we select the state $s^{*}$ that satisfies:
$s^{*}=\arg\max_{s_{i}\in\mathcal{S}}\left\{\sum_{j=0}^{i-1}-\log p(t_{j}|s_{j})\right\}$
where $\mathcal{S}$ represents the set of all unexpanded states, $(s_{0},t_{0}),\cdots,(s_{i-1},t_{i-1})$ denotes the proof trajectory leading to state $s_{i}$, and $\log p(t_{j}|s_{j})$ represents the average log probability of each generated token conditioned on the state. The algorithm expands up to $N$ states, achieving successful proof search upon reaching any proof state with no remaining goals. Multiple search attempts ($K$ times) can be performed to calculate a pass rate $pass@K$. However, analysis of the search trajectories of our models indicates that the best-first search method exhibits poor performance (as shown in Fig. 5). Thus, we hypothesize that average log probabilities may not serve as an effective metric to guide the heuristic search process.
Critic-Guided Search
Recent literature (Lin et al., 2024) indicates that best-first search (BFS) methods may be suboptimal for formal proving. A primary advantage of tree search strategies, when compared to whole-proof generation, is their capacity to leverage critical information from intermediate prover states. However, best-first search algorithms, as they only look at the average log probability of tactics, often fail to fully capitalize on this advantage, as shown in Fig 1. Therefore, using best-first-search with log-probability scores seldom leads to deep proofs and limits the proving ability of our model, which is consistent with our experiments (shown in Fig 3). Therefore, we choose to train a critic model (Lample et al., 2022; Polu et al., 2022) to better guide our prover model for proof generation.
The critic model ($V$) uses the proof state ($s$) as the input and outputs a scalar ($V(s)\in\mathbb{R}$). At each iteration, the algorithm selects the optimal state $s_{i}$ for expansion by querying the critic model. Then it employs the prover model to generate candidate tactics. We train our critic model in a preference style which is similar to reward model training in RLHF (Ouyang et al., 2022) instead of binary targets (the state can be proved or not). Two types of preference pairs are created in the training process:
  • Path Pairs: For a successful proof path from the root (initial proof state) to no_goals, we hypothesize that a state close to the target is always better. Thus we create preference pairs where the positive example is a child state closer to no_goals and the negative example is the parent state closer to the initial root state (i.e. $V(s_{t})<V(s_{t+\Delta})$), which implicitly assumes that any legal tactic on a successful path toward no_goals yields a positive reward $r$. For a path of length n, this methodology allows us to generate at most $n\choose 2$ pairs of positive and negative examples.
  • Sibling Pairs: We construct preference pairs consisting of a state on the successful path (positive example) and its sibling state (negative example) (i.e. $V(s_{sibling})<V(s_{t})$). Sibling states are defined as child nodes of the same parent that did not lead to no_goals. This design is based on the principle that the state on the successful path ($s_{t}$) is preferable to its sibling state ($s_{sibling}$), which did not lead to a proof.
2.2 Expert Iteration
Bootstrapping
The proposed approach utilizes the InternLM2-StepProver(Wu et al., 2024), our latest model for formal reasoning, as the foundational prover model, along with an initial dataset. This dataset is aggregated from four distinct sources: the miniF2F-train split, the mathlib dataset, the Lean-Workbook, and the Lean-Github dataset. Collectively, these sources furnish the initial states $s^{i}$, each representing a theorem to be proven. For the initialization of the critic model, the prover model is initially tasked with generating $K$ solution samples (trajectories) per theorem. Following this, a preference dataset is curated by sampling preference pairs from successful search trajectories; this dataset is then used for training the prover model. The overall iterative procedure is executed multiple times. The process terminates when the critic-guided search method achieves performance superior to that of standard (vanilla) methods on the miniF2F-train benchmark, which serves as an indicator that the critic model is stabilized.
Expert Iteration on Lean-Workbook
With our initialized prover and critic models, we perform expert iteration on the Lean-Workbook dataset. Due to limitations in auto-formalization accuracy, it is recognized that some formalized propositions might be incorrect. Therefore, every formalized proposition includes its negated version in the dataset, following the paradigm outlined in (Xin et al., 2024a). Proofs are searched using best-first and critic-guided algorithms by generating a tactic as an action (Polu et al., 2022; Azerbayev et al., 2023b; Wu et al., 2024).
@@ 2.2 Expert Iteration
Initially, we conduct a rapid scan of the entire Lean-workbook-plus dataset using a relatively small search budget (i.e., a maximum of 10 iterations per problem and a time limit of 50 seconds). The discovered proofs are added to the training set, and the solved problems, along with their negated statements, are removed from the dataset. This process helps us identify statements that are inherently unprovable, thereby enhancing the efficiency of the iteration.
@@
This process is then repeated over multiple rounds, gradually increasing the search budget for subsequent evaluations until a predefined upper bound (at most 2000 iterations and 3600 seconds per problem) is reached. After each round, we retrain our prover and critic models using an expanded set of successful proof trajectories. Since some found proofs are ill-formed and contain many irrelevant proof steps with higher CPU consumption, we continue to search for proofs for these problems, aiming to use shorter and more direct proofs to improve our models.
After model iterations, we use the critic model to re-estimate all unproven statements. Next iteration will then focus on the top 50% of problems that the critic model indicates are most likely to be solvable.
Prover Model Objective
@@
NAME: square_sub_one_divisible_eight
PROOF_BEFORE: rw [h, pow_two]
STATE_BEFORE: m n :
h : n = 2 * m + 1
8 | (2 * m + 1) * (2 * m + 1) - 1
TACTIC:
rw [← Nat.mod_add_div (2 * m + 1) 8]
Figure 2: An example of the prompt used by the prover model.
Figure 3: Critic-guided search finds more deep proofs compared to best-first-search in ProofNet and miniF2F-test. We calculate the proof length based on the number of tactics. Only the shortest 5 proofs are considered for each solved problem.
@@ 3 Experiments and Results
We evaluate our approach on Lean-workbook-plus (Ying et al., 2024a), one of the largest auto-formalized problem sets in Lean 4, which provides comprehensive coverage of diverse mathematical reasoning tasks. The integration of critic mechanisms enables our system to discover deeper proofs while maintaining computational efficiency. We trained InternLM2.5-StepProver and InternLM2.5-StepProver-critic on Lean-Workbook with more than 20000 CPU hours of expert iteration. The models are then evaluated on miniF2F(Zheng et al., 2021) and ProofNet(Azerbayev et al., 2023a).
Experimental results show that both critic-guided search and expert iteration boost the theorem proving capabilities of the prover model. The critic-guided search techniques significantly enhance the prover model in its ability to find deeper proofs. We describe our experiments and findings in detail below.
3.1 Benchmark Performance of InternLM2.5-StepProver
A comprehensive analysis of InternLM2.5-StepProver is conducted on several standard formal benchmarks, in comparison with our previous model InternLM2-StepProver, as well as a number of other frontier language models, to exhibit the strength of our approach.
Method Model size Pass miniF2F-valid miniF2F-test
Whole-Proof Generation Methods
TheoremLlama (Wang et al., 2024b) - cumulative $6.5\% 33$$3.6\%$
DeepSeek-Prover (Xin et al., 2024a) 7B 128128 - $46.1\%$
$16\times 4096$ - $50.0\%$
DeepSeek-Prover-V1.5-RL 7B 32 - $50.0\%$
64 - $50.7\%$
128128 - $51.6\%$
32003200 - $54.9\%$
$4\times 6400$ - $58.4\%$
$16\times 6400$ - $60.2\%$}
Tree Search Methods
PACT (Han et al., 2021) 837M $1\times 16\times 512$ $23.9\%$ $24.6\%$
$8\times 16\times 512$ $29.3\%$ $29.2\%$
ReProver (Yang et al., 2024) 229M - - $26.5\%$
Llemma (Azerbayev et al., 2023b) 7B $1\times 32\times 100$ $\% 26.226.2\%$
$8\times 8\times 512$ $41.2\%$ $34.5\%$
$64\times 8\times 512$ $47.3\%$ $36.6\%$
HTPS (Lample et al., 2022) 600M cumulative $58.6\%$ -
$64\times 5000$ - $41.0\%$
Lean-STaR  (Lin et al., 2024) 7B $64\times 1\times 50$ - $46.3\%$
InternLM2-Math-Plus 7B $1\times 32\times 100$ - $43.4\%$
DeepSeek-Prover-V1.5-RL 7B $1\times 3200$ - $55.0\%$
$4\times 6400$ - $59.6\%$
$16\times 6400$ - $62.7\%$}
$32\times 6400$ - $63.5\%$}
$64\times 32\times 100$ $63.9\%$} $54.5\%$}
InternLM2.5-StepProver-BF 7B $1\times 32\times 600$ $55.4\%$ $47.3\%$
$4\times 32\times 600$ $61.3\%$} $52.6\%$}
$16\times 32\times 600$ $63.7\%$} $57.3\%$}
$64\times 32\times 600$ $64.6\%$} $59.2\%$}
$256\times 32\times 600$ $65.1\%$} $59.4\%$}
InternLM2.5-StepProver-CG 7B $1\times 32\times 600$ $49.4\%$ $43.0\%$
$4\times 32\times 600$ $55.9\%$} $56.1\%$}
$16\times 32\times 600$ $64.5\%$} $61.7\%$}
$64\times 32\times 600$ $67.7\%$} $64.3\%$}
$256\times 32\times 600$ $68.4\%$} $65.6\%$}
InternLM2.5-StepProver-BF+CG 7B $2\times 32\times 600$ $56.0\%$ $50.7\%$
$4\times 32\times 600$ $61.4\%$} $58.5\%$}
$16\times 32\times 600$ $65.8\%$} $62.5\%$}
$64\times 32\times 600$ $68.0\%$} $63.8\%$}
$256\times 32\times 600$ 69.6% 65.9%
miniF2F
We first analyze the performance on the miniF2F benchmark (Zheng et al., 2021). The original benchmark was released in Lean 3 and was later ported to an earlier version of Lean 4. We use the Lean 4 version of miniF2F, as released by the LeanDojo project (Yang et al., 2024), with our adaptations to Lean 4.7.0 and corrections of several formalization mistakes.
The best-first-search approach employs an evaluation setting similar to that of InternLM2-StepProver, where the model selects states to expand based on the average log-likelihood of the tactics leading to those states. In contrast, the critic-guided (CG) search method involves the prover model selecting states to expand based on InternLM2.5-StepProver-critic that grades each state. The search budget for both methods can be universally described as $P\times S\times K$, where $P$ represents the number of passes, $S$ the number of states, and $K$ the maximum number of state expansions, or search iterations. In our context, we set $S=32$ and $K=600$, with the temperature fixed at $T=0.7$. In a BF+CG scenario, the computation budgets are equally distributed. The test results are presented in Tab. 1.
ProofNet
The evaluation setting is mainly the same as miniF2F. Tab. 2 presents the performance of various models on the ProofNet dataset. The BF+CG strategy equally distributes search budgets on two methods. As we are not taking the validation set of ProofNet for expert iteration, only the pass rate on the whole dataset is reported. InternLM2.5-StepProver achieved a pass@256 of 27.0% overall.
Critic-Guided Search Improves Theorem Proving.
From Tab. 1, it is observed that critic-guided search performs significantly better than the best-first method using the same prover model. The critic-guided method achieves an accuracy rate of 68.4% on miniF2F-Valid and 65.6% on miniF2F-Test, which is only slightly lower than the combined methods and substantially higher than the BF method. These experimental results indicate that the critic model enhances the inherent proving ability of the prover model, enabling it to discover deeper proofs. Analysis of the length distribution of the searched solutions (shown in Fig. 3) further validates this observation, as CG methods consistently discover deeper proofs than BF methods, successfully constructing proofs with more than 9 tactics.
Critic-Guided Search Explores a Distinct Proof Space.
Another observation is that the proof distribution discovered by critic-guided (CG) search differs significantly from that of the best-first (BF) approach. The synergistic effect of combining the two methods highlights this distinction. For instance, a hybrid model achieves 69.6% accuracy on miniF2F-Valid and 65.9% on miniF2F-Test. On the ProofNet benchmark, this synergy is even more pronounced: while BF and CG individually solve 22.3% and 23.9% of problems, their combination solves 27.0%. These results strongly indicate that critic-guided search navigates the proof space differently, identifying novel solutions that a scaled-up BF search alone would not discover. However, with low sampling budgets, CG may overlook simpler proofs that are trivial for BF. Therefore, a hybrid search was adopted to leverage the strengths of both methods in our experiment.
Table 2: Pass rates on the ProofNet (Azerbayev et al., 2023a) dataset.
Method Pass ProofNet
valid test all
ReProver (Yang et al., 2024) - - - $13.8\%$
Deepseek-Prover-V1.5-RL $1\times 3200$ $22.0\%$ $21.5\%$ $21.8\%$
$4\times 6400$ $25.4\%$ $\% 25.325.3\%$
InternLM2-StepProver $1\times 32\times 100$ - - $18.1\%$
InternLM2.5-StepProver-BF $2\times 32\times 600$ - - $18.8\%$
$32\times 32\times 600$ - - $21.1\%$
$128\times 32\times 600$ - - $22.3\%$
InternLM2.5-StepProver-CG $2\times 32\times 600$ - - $17.4\%$
$32\times 32\times 600$ - - $21.9\%$
$128\times 32\times 600$ - - $23.9\%$
InternLM2.5-StepProver-BF+CG $4\times 32\times 600$ - - $18.8\%$
$64\times 32\times 600$ - - $23.6\%$
$256\times 32\times 600$ - - $27.0\%$
@@ 3.2 Results of Expert Iteration
As Tab. LABEL:tab:leanwkbk_results shows, a total of 17.0% of the Lean-workbook-plus problems are proved or disproved, making a noticeable improvement since the release of Lean-Workbook. These proved and disproved statements and their corresponding tactics and states have been released. We also revealed more facts about the expert iteration process, especially the efficacy of our CG method.
@@
Scaling CPU/GPU Computation in Expert Iteration.
The search process involves a collaboration of CPU and GPU resources. Given the fixed amount of active GPUs and CPUs, the GPU time consumed is proportional to the total CPU time (in our case, approximately 1:11). In summary, approximately 21,364 CPU days are consumed throughout the entire expert iteration process. However, these search budgets are not uniformly distributed across all formalized problems. Easier problems are more likely to be solved in the early rounds of iteration, thereby ceasing to consume search resources in later rounds. The search consumption of each problem is a key indicator of the distribution of problem difficulty and can provide valuable insights for further scaling. In our case, we selected the CPU time consumed per successful proof as an estimate of resource consumption expectations. Consider the set of problems $S=\{s_{i}\}$. For each problem $s_{i}$, let $P_{s_{i}}$ denote the set of all attempts, and $T_{s_{i},j}$ represent the time spent on the $j$-th attempt of problem $s_{i}$. Define the indicator function $\text{valid}(s_{i},j)$, which equals 1 if the $j$-th attempt on problem $s_{i}$ results in a valid solution, and 0 otherwise. The CPU time consumed per successful proof, $C_{s_{i}}$, is given by: $C_{s_{i}}=\frac{\sum_{j}T_{s_{i},j}}{\sum_{j}\text{valid}(s_{i},j)}$.
Table 4 presents a detailed analysis of computational resource consumption. A key observation is that the majority of CPU resources are expended on problems that are challenging to prove or disprove. Notably, only about 1.5% of CPU resources are used to solve 17.0% of the problems, while the remaining 98.5% of resources yield no successful outcomes.
Fig 4 provides a more granular analysis of the distribution of CPU search time for each problem. A number of observations can be made from the graph. First, the graph reveals a peak in the near-zero region, suggesting the presence of numerous trivial problems in the auto-formalized dataset. Second, a log-linear trend is evident for problems with CPU search times between approximately 0 and 10,000 seconds. Assuming an equal distribution of problem difficulties, searching for longer proofs becomes exponentially harder due to the search space explosion problem, which explains the occurrence of this trend. Unfortunately, the best-first approach is observed to degrade to a brute-force approach—discovering fewer proofs with larger search budgets (¿10,000 s)—while the critic-guided approach performs better, discovering more solutions beyond the log-linear trend.
@@
Figure 4: The distribution of CPU search time per successful proof, which is defined as the total CPU resources spent on searching the specific problem divided by the number of independent successful search trials. Only problems solved are represented in the figure. 83% of the Lean-workbook problems that remain unproven are not depicted. BF and CG are analyzed independently.
Scaling Proof Lengths in Expert Iteration.
As analyzed in the previous section, best-first-search (BF) methods suffer from an explosion of the search space, rendering it inferior under large sampling budgets. In contrast, critic-guided (CG) methods are capable of finding longer solutions. The average length of solutions found by the BF method is 1.66, whereas the same indicator is 4.44 for the CG method, demonstrating a significant improvement in deeper reasoning abilities. The distribution of proofs during our expert iteration is shown in Fig. 5. To avoid the impact of redundancies, only the shortest proof for each problem is considered here. A nearly log-linear trend is observed in the distribution of proof lengths found by best-first search method, while the critic-guided proof method can find much longer proofs. However, 83% of the problems in the Lean-workbook remain unproven, which can have longer proof lengths and require much more time to find them based on this estimation. We hope that our proposed method can help facilitate the scaling of formal proof searching beyond the ”log-linear” boundaries.